# Two Heads Are Better Than One: A Multi-Agent System Has the Potential to Improve Scientific Idea Generation

### v1

@@ 4.1 Experimental Settings
Dataset We build our scientific research ecosystem using real scientists’ information from the AMiner Computer Science Dataset  [^1] https://www.aminer.cn/aminernetwork, which was constructed by extracting scientists’ profiles from online web databases (Tang et al., 2008). This dataset includes 1,712,433 authors and 2,092,356 papers, covering the period from 1948 to 2014, with disambiguated author names. To manage the large volume of data, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2000, 2010, and 2014, respectively. For quality assurance, we filtered out past papers lacking abstracts or with fewer than 10 citations, contemporary papers with fewer than 5 citations or missing abstracts, and authors with fewer than 50 papers or 50 co-authors. As a result, we extracted detailed information from 156 authors and 85,217 papers to construct the ecosystem and initialize the corresponding agents for the simulation. All paper and author data are embedded using the “mxbai-embed-large” model (Lee et al., 2024).
Implementation We implement our system on top of the Agentscope framework (Gao et al., 2024), which serves for LLM-empowered multi-agent applications. We evaluate our system using different publicly available LLMs: GPT-4o (OpenAI, 2023) and Llama-3.1 (8b and 70b) (Dubey et al., 2024). GPT-4o is accessible exclusively via a public API, while the Llama-3.1 models are open-weight and invoked using the Ollama (Ollama, 2024) in our experiments. Each experimental run on Llama-3.1 (8b) takes approximately 10 minutes on 1 NVIDIA A100 40G GPU within a team discussion setting of 4 members and 5 turns ($K$ = 5). All experimental results are averaged on 20 runs.
@@
Evaluation Metrics Since no single evaluation metric perfectly captures the novelty of scientific outputs, we employ three common metrics that align with our intuition: (1) Historical Dissimilarity (HD): The average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in $B_{\text{past}}$ (Shao et al., 2020; Zhou et al., 2024). A larger distance indicates that the generated abstract is more dissimilar from existing papers, suggesting a higher likelihood of novelty. (2) Contemporary Dissimilarity (CD): The average Euclidean distance between the generated abstract embedding and embeddings of the top 5 most similar abstracts in $B_{\text{con}}$. A smaller distance indicates greater similarity to newer papers, also suggesting a higher likelihood of novelty. (3) Contemporary Impact (CI): The average citation count of the top 5 most similar abstracts in $B_{\text{con}}$ (Yang et al., 2022). A higher citation count suggests that the generated abstract is more likely to have a higher impact. To ensure comparability, we normalize each calculated metric using the mean value derived from the entire corresponding database, with normalization defined as the metric divided by the mean value. Since novelty is difficult to measure directly, we introduce a proxy metric to comprehensively account for the three indicators: (4) Overall Novelty (ON). We assume that ON is positively related to both HD and CI and negatively related to CD, calculated as $\text{ON}=(\text{HD}\times\text{CI})/{\text{CD}}$. Mathematically, the expected value of ON is proportional to the true novelty. Further discussion of this metric is provided in the Appx. F.
Experimental Design We aim to explore several key aspects of VirSci’s performance. (1) How do the evaluation metrics of abstracts from VirSci outperform those generated by the SOTA single-agent method AI Scientist under similar conditions (Sec. 4.2)? (2) How do the novelty scores of abstracts improve in relation to different team settings, including team size and discussion turns (Sec. 4.3.1), team freshness (Sec. 4.3.2), and team research diversity (Sec. 4.3.3)? (3) How do different components in our system affect the novelty of the generated abstracts (Sec. 4.4)? (4) How do adaptive turn numbers affect the novelty of the generated abstracts (Sec.5)?
4.2 Comparisons with AI scientist
Table 1: Comparisons with AI Scientist. Results show that our multi-agent system outperforms the AI Scientist across all metrics, with GPT-4o achieving the highest performance.
Agent Model Method LLM Review Score $\uparrow$ CD $\downarrow$ CI $\uparrow$
GPT-4o1 AI Scientist 3.10 0.38 3.22
Ours 3.34 (+0.24) 0.34 (-0.04) 3.78 (+0.58)
LLaMA3.1-8b AI Scientist 2.09 0.49 2.12
Ours 2.31 (+0.22) 0.42 (-0.07) 3.29 (+1.17)
LLaMA3.1-70b AI Scientist 2.24 0.48 2.11
Ours 2.53 (+0.29) 0.40 (-0.08) 3.36 (+1.25)
  • 1
    GPT-4o API is “gpt-4o-2024-08-06”.
Adjusting Settings for Fair Comparison with AI Scientist Given that the problem formulation of the AI Scientist differs from ours, we must make several justifications to ensure relative fairness in our comparisons: (1) Since the AI Scientist is limited to generating ideas from its pre-defined topics (2D Diffusion, NanoGPT, and Grokking), we include NanoGPT in the topic selection prompt for VirSci as the initial discussion topic, ensuring that the final abstracts align with the same research direction. (2) Since our teamwork approach differs from the AI Scientist’s solitary investigation, we need to ensure that the comparisons are made under the same inference cost. The AI Scientist performs 50 turns of self-reflection during its idea generation, which does not apply to its paper generation. To align the inference costs, we set the number of team members to 4 and the number of discussion turns to 5, ensuring the experiments are conducted under approximately the same computational cost. (3) Since the AI Scientist lacks a scientific research ecosystem, it retrieves papers across all time ranges through the Semantic Scholar API (Fricke, 2018). To maintain consistency with the AI Scientist, we replace our databases with the Semantic Scholar API for both paper retrieval and metric calculation. Specifically, after generating ideas and corresponding abstracts from both the AI Scientist and our system, we use the generated ideas as queries to retrieve related papers, extracting corresponding abstracts and citation counts for evaluation. (4) We evaluate the generated abstracts using both our metrics (CD and CI) and the AI Scientist’s metric (LLM review score) (Lu et al., 2024). The LLM review score is calculated by GPT-4o which conducts abstract reviews based on a truncated version of the Neural Information Processing Systems (NeurIPS) conference review guidelines [^2] https://neurips.cc/Conferences/2024/ReviewerGuidelines, shown in Fig. 17.
Experimental Results and Analysis As shown in Fig. 1, our multi-agent system outperforms the AI Scientist across all three metrics: CD, CI, and LLM review score. This demonstrates that our approach effectively enhances the novelty of generated abstracts in a collaborative setting. Notably, GPT-4o consistently achieves the highest novelty scores among the models tested, reflecting its superior ability to generate innovative ideas and abstracts. In contrast, the Llama-3.1 (8b and 70b) models do not show a significant difference in novelty scores, suggesting that moderate changes in model size may not enhance novelty.
4.3 Exploring Science of Science: The Impact of Team Dynamics on Novelty
While the effects of team size, team freshness, and team research diversity on the novelty of research outputs have been established in the Science of Science field using traditional statistical methods (Wu et al., 2019; Zeng et al., 2021; Shi & Evans, 2023), they have yet to be verified in an LLM-based multi-agent system. We conduct the following experiments to demonstrate VirSci’s potential to simulate key findings in Science of Science.
4.3.1 How Team Size and Discussion Turns Affect Novelty
@@
Figure 3: Effects of team size and discussion turns on novelty scores. Peak innovation occurs with 8 members and 5 turns, while larger teams or excessive turns hinder creativity. “Inference Cost” is the product of team size and turns.
Effects of Team Size on Novelty The results shown in Fig. 3 indicate that increasing the number of team members can enhance the novelty score of the generated abstracts. By adding new team members, a broader range of ideas and perspectives can be facilitated, leading to more creative solutions and innovative outputs. However, this relationship is not strictly linear; our findings suggest that the peak novelty occurs with a team size of 8 members. While a moderate increase in team size may boost novelty, excessively large teams can introduce coordination challenges and communication barriers. These issues may dilute individual contributions and foster groupthink, where the focus shifts from original ideas to achieving consensus. Thus, there appears to be an optimal team size that maximizes creativity without overwhelming the collaborative process. This conclusion aligns with existing literature, which suggests that while smaller teams tend to disrupt science and technology with new ideas and opportunities, larger teams often concentrate on refining and developing existing concepts (Wuchty et al., 2007; Fortunato et al., 2018; Wu et al., 2019).
Effects of Discussion Turns on Novelty The number of discussion turns plays a crucial role in enhancing the novelty score (Mezirow, 2003; Li et al., 2023; Shinn et al., 2024; Lu et al., 2024). Our analysis (Fig. 3) indicates that an appropriate number of turns enables team members to explore topics thoroughly, iterate on ideas, and refine their abstracts. This iterative process is essential for deepening understanding and producing more sophisticated research outputs. While initial turns are beneficial for generating ideas and facilitating discussions, an excessive number of turns can lead to fatigue and reduced engagement. This may stifle creativity, as members may conform to dominant ideas rather than proposing innovative thoughts. Therefore, balancing the number of discussion turns is critical for maintaining high levels of novelty. Our findings suggest that peak novelty is achieved with a discussion turn count of 5.
Interaction Between Team Size and Discussion Turns Our findings further illustrate the interaction between team size and discussion turns. Larger teams with fewer turns can still produce relatively higher novelty scores, while smaller teams with excessive discussion turns may struggle to achieve the same level of innovation. This interplay emphasizes the importance of strategic planning in collaborative settings. Optimizing both team size and the number of discussion turns can significantly enhance the likelihood of generating novel outputs, indicating that the design of collaborative processes should take these factors into account.
4.3.2 How Team Freshness Affects Novelty
@@
Figure 4: The balance of new and returning collaborators in the team has a notable impact on novelty, with 50% freshness yielding the highest historical dissimilarity and overall novelty, particularly in larger teams (size 8).
As shown in Fig. 4, team freshness, the fraction of team members who have not previously collaborated, has a notable effect on the novelty of generated outputs. Notably, team freshness shows its strongest effects at 50%, particularly for larger teams (size 8). At this level, historical dissimilarity reaches its peak, suggesting that a balanced mix of new and returning collaborators promotes divergence from past research, enhancing overall innovation. As team freshness increases, contemporary dissimilarity decreases, indicating that teams with fresh members tend to generate abstracts that align more closely with future research trends. Furthermore, both CI and ON achieve their highest values at 50% freshness. This suggests that a balanced team composition, where half the members are new, optimally combines novelty and future relevance, driving impactful research outcomes. Although our findings focus specifically on novelty in scientific abstracts, the broader principle aligns, to some extent, with prior work in this field (Guimera et al., 2005; Zeng et al., 2021).
4.3.3 How Team Research Diversity Affect Novelty
@@
Figure 5: Effects of team diversity on novelty. The optimal diversity level appears to be 50%, which maximizes novelty and impact across team sizes.
@@
Team research diversity, defined as the proportion of team members who specialize in unique research topics different from others in the team, plays a significant role in influencing various novelty metrics. As shown in Fig. 5, increasing diversity enhances HD, with both team sizes showing a peak at 50% diversity, indicating that moderately diverse teams produce work that is distinct from previous research. CD decreases most notably at 50% diversity, suggesting that teams with balanced diversity align better with future research trends while maintaining innovation. In terms of contemporary impact, larger teams (size 8) benefit more from higher diversity, seeing a significant increase, while smaller teams exhibit a more stable, moderate impact. Finally, Overall Novelty is highest at 50% diversity, particularly for larger teams, reflecting the value of having a balanced mix of diverse and non-diverse members for producing novel and impactful work. This conclusion mirrors findings from the Science of Science, where an unexpected combination of team members can increase research impact (Uzzi et al., 2013; Shi & Evans, 2023).
4.4 Ablation Study
Effects of Components Designed for Novelty In this section, we respectively test the effects of the invitation mechanism in team discussion, the role of the novelty assessment step, and the impact of self-review in abstract generation. All experiments are conducted with a 5-turn discussion. The results consistently show improvements in ON when these components are applied. For the invitation mechanism (Tab. 4.4), introducing new scientists into the discussion positively impacts the team’s performance across both 4-member and 8-member teams. This indicates that seeking external insights from relevant but non-team scientists fosters more diverse and novel ideas. The novelty assessment step (Tab. 4.4) also significantly boosts the scores. If novelty assessment is not considered, then the output of idea generation will not be an idea list but the idea from the last scientist. Novelty assessment ensures that the generated ideas are continuously evaluated for originality, helping teams avoid overlap with existing research. The improvement is most noticeable in larger teams, where more ideas are being generated and assessed. Finally, the self-review mechanism (Tab. 4.4) is crucial in further refining the abstracts. By allowing the team leader to re-evaluate the abstract for novelty after it is fully generated, low-quality abstracts are discarded, and the team engages in a new discussion to generate a better idea, as evidenced by the score improvements for both team sizes.
Table 2: Effects of invitation mechanism.
Size Invitation ON $\uparrow$
4 - 3.30
3.40
8 - 4.12
@@ 4.4 Ablation Study
Table 3: Effects of novelty assessment.
Size Novelty Assessment ON $\uparrow$
4 - 3.19
3.40
8 - 3.98
4.23
Table 4: Effects of self-review in abstract generation.
Size Self-review ON $\uparrow$
4 - 3.26
3.40
8 - 3.99
4.23
Table 5: Comparison between fixed turns and adaptive turns in team discussions. The adaptive pattern shows both a lower inference cost and a higher ON.
Size Pattern Turns Inference Cost1 ON $\uparrow$

### v2

@@ 4.1 Experimental Settings
Datasets. We evaluate the performance of VirSci and baseline methods on two datasets: the AMiner Computer Science dataset [^5] https://www.aminer.cn/aminernetwork, containing 1,712,433 authors and 2,092,356 papers from 1948 to 2014, and the Open Academic Graph 3.1666https://open.aminer.cn/open/article?id=65bf053091c938e5025a31e2, comprising 35,774,510 authors and 130,710,733 papers as of 2023. For quality assurance, we applied filtering strategies, reducing the Computer Science dataset to 156 authors and 178,197 papers, and the Open Academic Graph to 3,169 authors and 201,131 papers. Additional dataset details and filtering strategies are provided in Appx. B.
@@
Evaluation Metrics.  In this paper, we adopt a more objective approach by using three common metrics that align with our intuition, as no single metric fully captures the novelty of scientific outputs: (1) Historical Dissimilarity (HD): The average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in $B_{\text{past}}$ (shao2020bert, ; zhou2024fine, ). A larger distance indicates greater dissimilarity from existing papers, suggesting a higher likelihood of novelty. (2) Contemporary Dissimilarity (CD): The average Euclidean distance between the generated abstract embedding and embeddings of the top 5 most similar abstracts in $B_{\text{con}}$. A smaller distance indicates greater similarity to newer papers, also suggesting a higher likelihood of novelty. (3) Contemporary Impact (CI): The average citation count of the top 5 most similar abstracts in $B_{\text{con}}$ (yang2022gender, ). A higher citation count suggests that the generated abstract is more likely to have a higher impact. To ensure comparability, we normalize each calculated metric using the mean value taken over all papers published in the same year as the chosen similar abstract in the corresponding database, with normalization defined as the metric divided by the mean value (alshebli2018preeminence, ).
Since novelty is difficult to measure directly, we introduce a proxy metric to account for the three indicators: (4) Overall Novelty (ON). It is natural to assume that ON is positively related to both HD and CI and negatively related to CD, calculated as $\text{ON}=(\text{HD}\times\text{CI})/{\text{CD}}$. Mathematically, the expected value of ON is proportional to the novelty.
Baselines. In this section, we compare VirSci with two recent agent-based systems for scientific discovery: the multi-agent system HypoGen qi2024large and the SOTA single-agent system AI Scientist lu2024ai . Given the differences in their pipelines, we adjust the experimental settings to a comparable level for a fair evaluation, elaborated in Appx. B.3. We then assess the results using our proposed metrics, LLM review scores, and human evaluations (Appx. B.4) to demonstrate the superiority of our method in idea generation.
Figure 3: Evaluation of abstracts using our overall novelty metric and human evaluation. The Pearson correlation coefficient of 0.52 indicates a positive correlation.
4.2 Analysis of Proposed Metrics
To assess the validity of our proposed overall novelty (ON) metric, we selected 200 abstracts generated from the Computer Science dataset, which were evaluated using both ON and by human experts in the field of computer science (details in Appx. B.4). For the human evaluations, we the novelty scoring framework from (si2024can, ) as a guideline. To ensure objectivity, independent researchers assessed each abstract, and the average score was calculated. As shown in Fig. 3, the Pearson correlation coefficient between ON and human ratings demonstrates a positive correlation, which supports the validity of our metric. We also explore the correlation between ON and LLM-based review (a common novelty measurement method (gu2024survey, )) in Appx. C and conduct further analysis.
Method LLM $\uparrow$ Proposed Metric Human Evaluation
CD $\downarrow$ CI $\uparrow$ Nov $\uparrow$ Fea $\uparrow$ Eff $\uparrow$
Agent Model: GPT-4o
HypoGen 3.02 0.36 3.10 4.78 4.24 4.43
AI Scientist 3.10 0.38 3.22 4.94 4.18 4.77
Ours 3.34 0.34 3.78 5.24 4.52 4.95
Agent Model: LLaMA3.1-8b
HypoGen 2.17 0.51 2.16 3.54 3.56 3.45
AI Scientist 2.09 0.49 2.12 3.66 3.52 3.63
Ours 2.31 0.42 3.29 4.08 3.74 3.69
Agent Model: LLaMA3.1-70b
HypoGen 2.18 0.49 2.13 3.57 3.61 3.52
AI Scientist 2.24 0.48 2.11 3.88 3.60 3.66
Ours 2.53 0.40 3.36 4.18 3.84 3.75
Table 1: Comparison with HypoGen and AI Scientist. The results demonstrate that our multi-agent system surpasses baseline models across all metrics. Here, “Nov”, “Fea”, and “Eff” denote human evaluation for “Novelty”, “Feasibility”, and “Effectiveness”, respectively.
(a) Experiments on the Computer Science dataset.
(b) Experiments on the Open Academic Graph dataset.
@@ 4.2 Analysis of Proposed Metrics
Figure 4: Effects of team size and discussion turn on novelty. Peak occurs with 8 members and 5 turns, while larger teams or excessive turns hinder creativity. “Inference Cost” is the product of team size and turns.
(a) Experiments on the Computer Science dataset.
(b) Experiments on the Open Academic Graph dataset.
@@
Figure 5: The balance of new and returning collaborators in the team has a notable impact on novelty, with 50% freshness yielding the highest historical dissimilarity and overall novelty, particularly in larger teams.
(a) Experiments on the Computer Science dataset.
(b) Experiments on the Open Academic Graph dataset.
@@
Figure 6: Effects of team diversity on novelty. The optimal diversity level appears to be 50% or 75%, which maximizes novelty and impact across team sizes.
4.3 Comparisons with Baselines
As shown in Tab. 1, our multi-agent system outperforms baseline models across metrics: our proposed metrics (CD, CI), LLM review, and human evaluation (Novelty, Feasibility, and Effectiveness), demonstrating its effectiveness in enhancing abstract novelty through collaboration. Notably, GPT-4o-based agents consistently achieve the highest novelty scores, reflecting its superior idea generation capability. In contrast, LLaMA3.1-8b and LLaMA3.1-70b-based agents show similar novelty scores, suggesting that moderate model size changes may not improve novelty.
4.4 Exploring Collaboration Mechanism
While the dynamics of group collaboration have been extensively studied in the Science of Science, its applicability to artificial multi-agent systems and the effects it may produce remain unclear. In this section, we analyze factors influencing idea novelty in our system, including team size, freshness, and research diversity—factors previously analyzed in human teams (wu2019large, ; zeng2021fresh, ; shi2023surprising, ).
Effects of Team Size on Novelty. The results in Fig. 4 show that increasing team size can enhance the ON of generated abstracts by bringing in diverse ideas and perspectives. However, this relationship is not strictly linear, with peak ON occurring at a team size of 8. While moderate team expansion boosts novelty, excessively large teams can create coordination and communication challenges, leading to diluted contributions and groupthink. This aligns with existing literature, which suggests that smaller teams generate more disruptive ideas, while larger teams focus on refining existing concepts(wuchty2007increasing, ; fortunato2018science, ; wu2019large, ).
Effects of Discussion Turn on Novelty.  The number of discussion turns plays a crucial role in fostering novel ideas (mezirow2003critical, ; li2023camel, ; shinn2024reflexion, ; lu2024ai, ). As shown in our experimental results (Fig. 4), an optimal number of turns promotes deeper insights and more innovative research outputs. While the initial turns facilitate idea generation, excessive turns can result in fatigue and reduced engagement, ultimately hindering creativity. Our findings suggest that peak ON is achieved at around 5 discussion turns.
Consistency Between Two Datasets. As shown in Fig. 4, no significant differences were found in outputs from single-agent teams (team size equals 1) across the two datasets, as individual research lacks interdisciplinary input. The overall trend—novelty rising, peaking at a team size of 8 and 5 turns, then declining—was consistent across both datasets, demonstrating our platform’s robustness and the potential of multi-agent systems in enhancing scientific idea generation. These findings also support the value of interdisciplinary collaborations for driving higher-impact research (shi2023surprising, ).
Effects of Team Freshness on Novelty. As shown in Fig. 5, team freshness—the fraction of members who have not previously collaborated—affects the novelty of outputs. Freshness has its strongest effect at 50%, especially for larger teams (size 8), where HD peaks. This suggests that a balanced mix of new and returning members promotes innovation by diverging from past research. As freshness increases, CD decreases, indicating that teams with more fresh members align their ideas more closely with future research trends. Both CI and ON reach their highest values around 50% freshness, highlighting that a balanced team optimally combines novelty and future relevance, consistent with previous work (guimera2005team, ; zeng2021fresh, ).
@@ 4.4 Exploring Collaboration Mechanism
Effects of Team Research Diversity on Novelty. Team research diversity is defined as the proportion of team members specializing in distinct research topics. As shown in Fig. 6, moderate diversity boosts HD, with peak performance at 25-50% diversity for 4-member teams and 50-75% for 8-member teams across both datasets. CD drops before reaching 50% diversity, suggesting that diverse teams align better with emerging research trends while maintaining innovation. Larger teams benefit more from higher diversity, showing an increase in CI, while smaller teams exhibit more stable, moderate effects. Overall, ON follows a reverse U-shaped curve for both team sizes and datasets, underscoring the importance of balancing research diversity for novel outcomes. This conclusion mirrors findings in Science of Science, where unexpected team combinations can enhance research impact (uzzi2013atypical, ; shi2023surprising, ).
5 Ablation Study
5.1 Effects of Paper Database
Database ON $\uparrow$
Idea Generation Novelty Assessment
- - 2.60
- 3.62
2.76
@@ 5.1 Effects of Paper Database
Table 2: Effects of the paper database in the processes of idea generation and novelty assessment on the Computer Science dataset.
Database ON $\uparrow$
Idea Generation Novelty Assessment
- - 2.41
- 3.99
2.60
4.65
Table 3: Effects of the paper database in the processes of idea generation and novelty assessment on the Open Academic Graph dataset.
In the workflow of VirSci, the past paper database is utilized as a reference during idea generation and novelty assessment. To verify the role of the constructed paper database, we conduct ablation studies on both datasets under the setting of 8 team members and 5 discussion turns, as shown in Tab. 2 and 3. The experimental results demonstrate that having references enhances the novelty of the final generated ideas, avoiding shallow or fanciful outcomes produced by the system.
5.2 Effects of Components Designed for Improving Novelty
In this section, we respectively test the effects of the invitation mechanism in team discussion, the role of the novelty assessment step, and the impact of self-review in abstract generation. All experiments are conducted with a 5-turn discussion. The results consistently show improvements in ON when these components are applied.
Team Size Invitation Turns
1 2 3 4 5 6 7
- 1.67 0.75 2.29 1.92 3.30 2.78 2.47
4
1.75 0.90 2.36 2.00 3.40 2.91 2.53
- 3.36 3.53 3.88 3.49 4.12 3.37 3.30
8
3.48 3.67 3.97 3.56 4.23 3.48 3.43
Table 4: Effects of invitation mechanism in team discussion. Comparison results show that this mechanism helps to improve the novelty of outputs. The evaluation metric is ON.
Team Size Assessment Turns
1 2 3 4 5 6 7
- 1.56 0.73 2.20 1.87 3.19 2.74 2.39
4
1.75 0.90 2.36 2.00 3.40 2.91 2.53
- 3.28 3.41 3.82 3.43 3.98 3.27 3.25
8
3.48 3.67 3.97 3.56 4.23 3.48 3.43
Table 5: Effects of novelty assessment. Comparison results show that novelty assessment helps to improve the novelty of outputs. The evaluation metric is ON.
5.2.1 Invitation Mechanism
For the invitation mechanism (results are presented in Tab. 4), introducing new scientists into the discussion positively impacts the team’s performance across both 4-member and 8-member teams. This indicates that seeking external insights from relevant but non-team scientists fosters more diverse and novel ideas.
5.2.2 Novelty Assessment Mechanism
The novelty assessment step (results are presented in Tab. 5) also influences the scores. If novelty assessment is not considered, then the output of idea generation will not be an idea list but the idea from the last scientist. Novelty assessment ensures that the generated ideas are continuously evaluated for originality, helping teams avoid overlap with existing research. The improvement is more noticeable in larger teams, where more ideas are being generated and assessed.
Team Size Self-review Turns
1 2 3 4 5 6 7
- 1.60 0.74 2.26 1.89 3.26 2.77 2.41
4
1.75 0.90 2.36 2.00 3.40 2.91 2.53
- 3.32 3.44 3.84 3.45 3.99 3.33 3.27
8
3.48 3.67 3.97 3.56 4.23 3.48 3.43
Table 6: Effects of self-review in abstract generation. Comparison results show that self-review after abstract generation helps to check and improve the novelty of outputs. The evaluation metric is ON.
5.2.3 Self-review Mechanism
Finally, the self-review mechanism (results are presented in Tab. 6) is crucial in further refining the abstracts. By allowing the team leader to re-evaluate the abstract for novelty after it is fully generated, low-quality abstracts are discarded, and the team engages in a new discussion to generate a better idea, as evidenced by the score improvements for both team sizes.
5.3 Effects of Discussion Pattern
5.3.1 Discussion Topologies
Figure 7: Two different discussion topologies: sequential mode (Left), and random mode (Right). In sequential mode, participants take turns presenting their ideas in a round-table format, one at a time. In random mode, after an agent speaks, the next speaker is chosen randomly, with the restriction that the previous speaker cannot be selected consecutively. Note that the actual discussion topology of our system is more complex, benefiting from the proposed invitation mechanism.
Team Size Topology Turns
1 2 3 4 5 6 7
4 Sequential 1.75 0.90 2.36 2.00 3.40 2.91 2.53
Random 1.67 0.85 2.34 2.01 3.35 2.84 2.47
8 Sequential 3.48 3.67 3.97 3.56 4.23 3.48 3.43
Random 3.50 3.63 3.88 3.52 4.16 3.40 3.39
Table 7: Comparison between two discussion topologies: sequential mode and random mode. The results show that sequential mode generally outperforms random mode in most cases as it leverages more background knowledge.
To further explore the effect of the discussion pattern, particularly discussion topologies, we compare two modes: sequential mode and random mode. As shown in Fig. 7, in sequential mode, participants take turns presenting their ideas in a round-table format; in random mode, after an agent speaks, the next speaker is chosen randomly, with the restriction that the previous speaker cannot be selected consecutively. Note that the actual discussion topology of our system is more complex, benefiting from the proposed invitation mechanism. The comparison results are presented in Tab. 7, with tests conducted for team sizes of 4 and 8. It can be observed that sequential mode generally outperforms random mode in most cases, as it leverages a broader knowledge base from the scientist agents. In random mode, some agents remain silent (i.e., are skipped) when the number of turns is limited. As a result, sequential mode is selected as the default discussion topology.
5.3.2 Discussion Turns
[…]